# Integrating New Technologies into Science:

### 15 Dec 2023 v1

@@ Abstract
New technologies have the power to revolutionize science. It has happened in the past and is happening again with the emergence of new computational tools, such as Artificial Intelligence (AI) and Machine Learning (ML). Despite the documented impact of these technologies, there remains a significant gap in understanding the process of their adoption within the scientific community. In this paper, we draw on theories of scientific and technical human capital (STHC) to study the integration of AI in scientific research, focusing on the human capital of scientists and the external resources available within their network of collaborators and institutions. We validate our hypotheses on a large sample of publications from OpenAlex, covering all sciences from 1980 to 2020. We find that the diffusion of AI is strongly driven by social mechanisms that organize the deployment and creation of human capital that complements the technology. Our results suggest that AI is pioneered by domain scientists with a ‘taste for exploration’ and who are embedded in a network rich of computer scientists, experienced AI scientists and early-career researchers; they also come from institutions with high citation impact and a relatively strong publication history on AI. The pattern is similar across scientific disciplines, the exception being access to high-performance computing (HPC), which is important in chemistry and the medical sciences but less so in other fields. Once AI is integrated into research, most adoption factors continue to influence its subsequent reuse. Implications for the organization and management of science in the evolving era of AI-driven discovery are discussed.
  • The subscription opened at Baltimore, for this end extended thence to all the world – Urbi et orbi.
    Jules Verne, From the Earth to the Moon
@@ 1 Introduction
The pace of scientific progress is a direct correlate of our alliance with new technologies. It has been so in the past with instruments such as the microscope and telescope, more recently with computers and the internet (Rosenberg, 1992; Ding et al., 2010). Today the spotlight is on AI [^2] For simplicity, throughout the article, we will use the appellation “AI” instead of “AI/ML”., which is fast emerging in the scientific landscape as a tool with great potential for discovery, a new “general method of invention”, as many have recently argued (Agrawal et al., 2018; Cockburn et al., 2018; Furman and Teodoridis, 2020; Bianchini et al., 2022). Problems that once were impossible to contemplate, or even formulate, come around every day. Examples include predicting the 3D structure of proteins (Jumper et al., 2021), regulating nuclear fusion plasma in the tokamak configuration (Degrave et al., 2022), predicting the formation of the structure of the universe (He et al., 2019), and creating a map of the brains of small insects (Winding et al., 2023). The potential of AI to accelerate and advance scientific discovery is being explored across almost all scientific disciplines, and at different stages of research processes (see, e.g., Wang et al., 2023).
@@
A recent Nature survey (Van Noorden and Perkel, 2023) of more than 1,600 researchers worldwide asked scientists who have already experimented with AI what obstacles prevent them from using it more and received the following responses: ‘Lack of skills or skilled researchers’ (80%), ‘Lack of training resource’ (70%), ‘Lack of funding’ (50%), Lack of computing resources (35%), ‘Lack of data’ (30%), and others (18%). In contrast, scientists who have no experience with AI mostly declared that they do not find it useful for their research. It thus seems clear that AI adoption is inextricably linked, at a very least, to an insufficient understanding of the technology’s potential in a specific domain, individual skills, and the accessibility of computing capacity and data (also confirmed in, e.g., Hwang, 2018; Thompson et al., 2022). The bibliographic study by Thu et al. (2022), on the composition of scientific teams in applied AI, shows that the lack of skills on the part of domain scientists is often compensated for by engagement in interdisciplinary research with computer scientists, and that such interdisciplinary efforts are also successful in terms of citations received. However, not all scientists may be in a position to implement, or even consider, such a collaborative strategy. What (other) resources then favor AI adoption? Also, not all scientists who try AI intend to adopt it persistently in their research – in our sample that share is about 50%. What are the barriers that prevent these scientists from integrating AI into their research?
@@
Scientific knowledge and expertise will be measured on past publications, assessing thematic diversity, impact, collaborations with peers and other dimensions; the institutional setting will encompass the quality of the researcher’s home institution, reputation, research orientation, and access to high-performance computing (HPC), presumably crucial for modern AI research. As for the econometric strategy, we will model AI adoption in research papers as a function of technology available in a given field at a particular time and various aspects of scientists’ STHC. Our approach, based on conditional logit regression with matched pairs (adopters vis-à-vis non-adopters), allows us to remove other unobservable components – i.e. cohort and field interactions with dynamic AI technologies – that might influence the adoption decision.
@@ 2.1.1 Social capital and about-knowledge from social ties
About-knowledge can be thought as the kind of ‘know-how’ that is necessary to achieve an early point of connection with another field – “[A] range of fairly simple facts and information about the sort of problem domains and approaches that populate different fields and specialisms(Priaulx and Weinel, 2018, p. 8). It is not an intimate understanding of how the scientist’s core domain expertise interacts with another field, but rather a set of cognitive foundations into the kind of contributions that the field can make to their own expertise, as well as a broad understanding of the practical settings, languages, sub-cultures, expectations and reward models that regulate that field. [^3] It is noteworthy that this type of knowledge and resulting connections, which we believe can be immensely valuable, particularly in the initial stages of the collaborative life cycle, are often overlooked in the literature on interdisciplinarity and transdisciplinarity.
It is clear that the lack of insights into what AI scientists do and/or where AI research is headed significantly decreases the likelihood that a scientist will recognize the relevance of the field’s contribution to their own work. In this sense, about-knowledge should fill the deficit by providing a wide-angled lens of the potential of the technology, thus prompting the scientist to give it a try.
@@
So let us start from collaboration networks. Here, we use the network of past collaborations of domain scientists as a valuable historical record of their interactions and collaborations with peers (see, e.g., Katz and Martin, 1997), particularly those with some experience in AI technology within the same application domain or in computer science more in general. It is reasonable to assume, in fact, that scientists who are embedded in a network where their peers have already proven experience with a certain technology (e.g., as evidenced by at least one publication) have more incentives to adopt the same technology in their own research, especially if their colleagues have achieved successful outcomes. This is because, even if a collaboration does not occur with a past collaborator, being part of a network reinforces a scientist’s rudimentary knowledge about the potential of the technology. More formally, we argue that:
@@ 2.1.2 Mentorship and newbies
The academic job market is rich in human resources specializing in AI. According to some recent statistics, the number of AI-related curricula has increased more than any other curriculum in recent years and is unlikely to slow down in the years to come. For instance, in 2020 alone, over 30,000 undergraduate students in the US completed a computer science degree, and one in every five students who earned a PhD degree in computer science specialized in AI  [^5] See Stanford AI Index Report from 2022 here: https://aiindex.stanford.edu/report/. We can expect similar figures in many other countries. AI-related courses are no longer limited to computer science departments at the undergraduate level; rather, a growing number of universities offer interdisciplinary programs that combine AI with other fields. The new generation of scientists also has at their disposal a plethora of online resources offered by universities and private companies that focus specifically on AI. One example is Massive Open Online Courses (MOOCs), which are emerging as an affordable and popular option for those who want to deepen their knowledge of AI, from introductory courses to others on cutting-edge algorithms and advanced applications. [^6] The importance of AI literacy from the early stage of education has also been recognized globally. A recent report by UNESCO (2022) highlights the commitment of several countries to developing AI literacy and competencies in K-12 schools. Generally, these initiatives aim to prepare new generations for a world in which AI will be ubiquitous, and thus understand the power and versatility of this technology along with its ethical dilemmas.
@@ 2.1.3 Computational resources
While AI is commonly perceived as an intangible technical system, it is de facto rooted in physical infrastructure and hardware. Yet, the role of physical assets and their associated computing capabilities – also known as AI compute – have been largely overlooked in policy circles and scholarly literature. [^7] One reason for that is the lack of standardized and validated data on computing resources. National and institutional and data on the supply and demand of AI compute is not easily accessible and, in some cases, considered sensitive proprietary information.
@@
Here, we measure the availability of AI compute by the presence of an HPC cluster within the focal researcher’s organization (although it should be noted that such compute infrastructure can also be used for non-AI workloads such as mathematical modeling and simulations). In very general terms, HPC is a technology that uses clusters of powerful processors, working in parallel, to process data and solve complex problems at high speeds (OECD, 2023). Unlike standard computing systems, HPC systems can handle multiple tasks simultaneously across multiple computer servers or processors with a centralized scheduler that manages the computing workload. The high cost of HPC can put this technology out of reach for most organizations, resulting in a significant “compute divide” within and between countries and institutions, as well as between the private sector and academia (Ahmed and Wahed, 2020; Ahmed et al., 2023). This is especially true for AI applications in some data-intensive scientific fields such as genomics and bioinformatics or particle physics where ML training and inferencing can be highly demanding in terms of memory and computational resources. The existence of a computational divide can therefore impede the adoption of AI and generate disparities in the productivity gains that AI can offer to science. In summary, our hypothesis is that:
@@ 2.2.2 Taste for exploration
Exploration is intimately linked to curiosity, a personal trait that prompts individuals to explore uncharted territories. We think that curiosity is a useful construct for understanding scientists’ behavior in terms of technology adoption. Although psychologists have not reached a consensus on its definition, it is generally accepted that curiosity involves an intrinsic motivated desire for new information – an “appetite for knowledge”, or more formally “a form of cognitively induced deprivation that arises from the perception of a gap in knowledge or understanding(Loewenstein, 1994, p. 75). [^8] It could be argued that an individual may also be curious about the topics she knows best. However, it should be noted that our definition of curiosity extends beyond the inclination of a scientist to expand her understanding within the area she is most knowledgeable about – a characteristic that should be common to every scientist! – but rather encompasses the search for knowledge and information far out in the knowledge space.
@@ 3.1 Data
We use the entire OpenAlex database, which had more than 230 million scientific articles as of August 2022. OpenAlex provides us with information relevant to our study, namely titles and abstracts. Authors are disambiguated so that we can track the paper trail of our focal scientists and their co-authors. Thus, we can measure the evolution of the network of co-authors and the (publication) experience of the scientists who are part of it. Author affiliations are also clean and geographically localized. We add additional information about the organizations in our sample, particularly the university’s Shanghai ranking and the availability of high-performance computing (more on this below). Finally, we rely on a scientific journal categorization system provided by OpenAlex, called ‘concept’. In particular, we denote the scientific field of a paper by 0-level concept assigned to the journal of the paper. [^10] There are 19 concepts at level zero, and five layers of concepts derived from them, representing a total of around 65,000 concepts. Level zero concepts include Medicine, Physics, Biology, Chemistry, etc.
@@
The development of AI as a research tool is a relatively recent occurrence, and the widespread adoption of deep learning applications gained momentum in the early 2010s (Bianchini et al., 2022). More specifically, 2012 can be considered the beginning of the AI revolution, with significant advances in neural network technology (Krizhevsky et al., 2017; Sevilla et al., 2022). This prompts us to restrict attention to the first AI use in the period from 2012 (the year when AI took off) to 2020 (end of sample period). We further limit the analysis to researchers who started research after 1980; excluding older scientists at the end of their careers and, hence, in general less susceptible to adopt AI in research. [^11] Researchers who began before 1980 are approaching the end of their careers, and thus their AI adoption dynamics may be less influenced by their scientific and technical human capital, and more affected by impending retirement. Our sampling definition of focal scientists is concisely described in Figure 1.
@@ 3.3 Econometric strategy
where $p(y_{i,t}=1)$ is the probability that a paper produced applies AI conditional on a paper produced at all, $\phi$ denotes the logit-function. $\mathbf{h}_{i,t-1}$ is the log-transform of our measurements $\mathbf{H}_{i,t-1}$. In principle, some relevant capital and/or individual specific tendencies may be also unobserved $\nu_{i,t}$. [^12] For convenience, the coefficients associated with AI research technology and STHC have the same names in the paper production function and the adoption function, but they are of course not the same. With this estimation equation, the dynamics of AI technologies is effectively dealt with by introducing intercepts $\gamma_{s,t}$ that vary over time and scientific domains.
@@
These common factors can be removed in a conditional logit approach. First note that matching is on the outcome such that in all matches one individual $i$ adopts AI and the other scientist $j$ does not adopt AI $(y_{i,t}=1,y_{j,t}=0)$, or vice versa. The conditional logit model, takes into account that only two possible outcomes are possible, and we estimate the probability of one of them:
@@
where common factors have been pulled out of the second equation.
@@ 4.1 Main trends on AI adoption in science
A growing number of players: The adoption of AI technology by domain scientists is on the rise, with over 20,000 individuals incorporating it for the first time in 2020 alone. This pattern aligns with recent research on the diffusion of AI in science, showing that the number of AI publications has surged nearly five-fold in the past decade, constituting around 5% of total scientific publications. Most of this research has gradually shifted from AI development to its application, which currently represents some 70% of scientific activity (Arranz et al., 2023). Surprisingly, despite technological advancements and increased resources for AI, the percentage of first-time users reusing technology remains low and stable at around 54% (13,211 re-use out of 23,918 first-use).
@@ 4.2.1 First-time AI adoption
Taken at face value the estimated coefficients suggest that being affiliated with an institution specialized in AI (AI inst. spe.=1), rather than not (AI inst. spe.=0), increases the odds-ratio for the focal scientist adopting AI compared to the matched scientist by 80%. [^13] To appreciate this point, note that our conditional logit model estimated on matched pairs allows for exactly two outcomes: either the focal scientist adopts AI and the matched scientist does not, or vice versa. The odds-ratio of the focal scientist adopting AI over the matched scientist adopting AI is $exp(\beta h_{i})/exp(\beta h_{j})$. A change of $i$’s factor from $h_{i}$ to $h^{\prime}_{i}$ thus implies a relative change of the odds-ratio by $exp(\beta(h^{\prime}_{i}-h_{i})$. A binary factor changing from 0 to 1 yields a relative change of $exp(\beta)$, and a continuous factor (in logs in the model) being higher by 10% is about $exp(\beta\times 1.1)$. Hence, interpretation of coefficient estimates is easily facilitated by considering the exponential of the coefficient estimate. Scientists of more ‘excellent’ institutions with 10% higher citation rates have 30% higher odds-ratios of AI adoption.

### 22 Feb 2024 v2

@@ Abstract
New technologies have the power to revolutionize science. It has happened in the past and is happening again with the emergence of new computational tools, such as artificial intelligence and machine learning. Despite the documented impact of these technologies, there remains a significant gap in understanding the process of their adoption within the scientific community. In this paper, we draw on theories of scientific and technical human capital to study the integration of AI in scientific research, focusing on the human capital of scientists and the external resources available within their network of collaborators and institutions. We validate our hypotheses on a large sample of publications from OpenAlex, covering all sciences from 1980 to 2020, and identify a set key drivers and inhibitors of AI adoption and use in science. Our results suggest that AI is pioneered by domain scientists with a ‘taste for exploration’ and who are embedded in a network rich of computer scientists, experienced AI scientists and early-career researchers; they come from institutions with high citation impact and a relatively strong publication history on AI. The access to computing resources only matters for a few scientific disciplines, such as chemistry and medical sciences. Once AI is integrated into research, most adoption factors continue to influence its subsequent reuse. Implications for the organization and management of science in the evolving era of AI-driven discovery are discussed.
@@ 1 Introduction
The pace of scientific progress is a direct correlate of our alliance with new technologies. It has been so in the past with instruments such as the microscope and telescope, more recently with computers and the internet (Rosenberg, 1992; Ding et al., 2010). Today the spotlight is on AI/ML, which is fast emerging in the scientific landscape as a tool with great potential for discovery, a new “general method of invention”, as many have recently argued (Agrawal et al., 2018; Cockburn et al., 2018; Furman and Teodoridis, 2020; Bianchini et al., 2022; Hain et al., 2023). [^2] For simplicity, throughout the article, we will use the appellation “AI” instead of “AI/ML”. Problems that once were impossible to contemplate, or even formulate, come around every day. Examples include predicting the 3D structure of proteins (Jumper et al., 2021), regulating nuclear fusion plasma in the tokamak configuration (Degrave et al., 2022), predicting the formation of the structure of the universe (He et al., 2019), and creating a map of the brains of small insects (Winding et al., 2023). The potential of AI to accelerate and advance scientific discovery is being explored across almost all scientific disciplines, and at different stages of research processes (see, e.g., Wang et al., 2023).
@@
A recent Nature survey (Van Noorden and Perkel, 2023) of more than 1,600 researchers worldwide asked scientists who have already experimented with AI what obstacles prevent them from using it more and received the following responses: ‘Lack of skills or skilled researchers’ (80%), ‘Lack of training resource’ (70%), ‘Lack of funding’ (50%), Lack of computing resources (35%), ‘Lack of data’ (30%), and others (18%). In contrast, scientists who have no experience with AI mostly declared that they do not find it useful for their research. It thus seems clear that AI adoption is inextricably linked, at a very least, to an insufficient understanding of the technology’s potential in a specific domain, individual skills, and the accessibility of computing capacity and data (also confirmed in, e.g., Hwang, 2018; Thompson et al., 2022). The bibliographic study by Thu et al. (2022), on the composition of scientific teams in applied AI, shows that the lack of skills on the part of domain scientists is often compensated for by engagement in interdisciplinary research with computer scientists, and that such interdisciplinary efforts are also successful in terms of citations received. However, not all scientists may be in a position to implement, or even consider, such a collaborative strategy. What (other) resources then favor AI adoption? Also, not all scientists who try AI intend to adopt it persistently in their research – in our sample that share is about 50%. What are the barriers that prevent these scientists from integrating AI into their research? [^3] The study of AI diffusion and, especially, adoption has hitherto focused on firms and not, to our knowledge, on scientists. In recent years, research has examined the various technological, organizational and environmental prerequisites for the (successful) adoption and use of AI technologies in companies, considering factors such as company size, industry sector and digital skills (see, e.g., Bughin et al., 2017; Alsheibani et al., 2018; Enholm et al., 2022; Kinkel et al., 2022; McElheran et al., 2024). This strand of literature has been helpful for our conceptual framework because several factors important to business find close parallels in the context of science – e.g., the size of a research team or the peculiarities of a scientific domain.
@@
Scientific knowledge and expertise will be measured on past publications, assessing thematic diversity, impact, collaborations with peers and other dimensions; the institutional setting will encompass the quality of the researcher’s home institution, reputation, research orientation, and access to high-performance computing (HPC), presumably crucial for modern AI research. As for the econometric strategy, we will model AI adoption in research papers as a function of technology available in a given field at a particular time and various aspects of scientists’ STHC. Our approach, based on conditional logit regression with matched pairs (adopters vis-à-vis non-adopters), allows us to remove other unobservable components – i.e., cohort and field interactions with dynamic AI technologies – that might influence the adoption decision.
@@ 2.1.1 Social capital and about-knowledge from social ties
About-knowledge can be thought as the kind of ‘know-how’ that is necessary to achieve an early point of connection with another field – “[A] range of fairly simple facts and information about the sort of problem domains and approaches that populate different fields and specialisms(Priaulx and Weinel, 2018, p. 8). It is not an intimate understanding of how the scientist’s core domain expertise interacts with another field, but rather a set of cognitive foundations into the kind of contributions that the field can make to their own expertise, as well as a broad understanding of the practical settings, languages, sub-cultures, expectations and reward models that regulate that field.
It is clear that the lack of insights into what AI scientists do and/or where AI research is headed significantly decreases the likelihood that a scientist will recognize the relevance of the field’s contribution to their own work. Recent research has shown that the same dynamics apply to AI adoption and diffusion within companies: uncertainty about the technology’s use cases (i.e., what to do with it and what competitive advantages it can bring) slows down adoption (Ameye et al., 2023). In this sense, about-knowledge should fill the deficit by providing a wide-angled lens of the potential of the technology, thus prompting the scientist to give it a try.
@@
So let us start from collaboration networks. Here, we use the network of past collaborations of domain scientists as a valuable historical record of their interactions and collaborations with peers (see, e.g., Katz and Martin, 1997; Newman, 2001; Wang et al., 2017), particularly those with some experience in AI technology within the same application domain or in computer science more in general. It is reasonable to assume, in fact, that scientists who are embedded in a network where their peers have already proven experience with a certain technology (e.g., as evidenced by at least one publication) have more incentives to adopt the same technology in their own research, especially if their colleagues have achieved successful outcomes. This is because, even if a collaboration does not occur with a past collaborator, being part of a network reinforces a scientist’s rudimentary knowledge about the potential of the technology. More formally, we argue that:
@@ 2.1.2 Mentorship and newbies
The academic job market is rich in human resources specializing in AI. According to some recent statistics, the number of AI-related curricula has increased more than any other curriculum in recent years and is unlikely to slow down in the years to come. For instance, in 2020 alone, over 30,000 undergraduate students in the US completed a computer science degree, and one in every five students who earned a PhD degree in computer science specialized in AI. [^5] See Stanford AI Index Report from 2022 here: https://aiindex.stanford.edu/report/ We can expect similar figures in many other countries. AI-related courses are no longer limited to computer science departments at the undergraduate level; rather, a growing number of universities offer interdisciplinary programs that combine AI with other fields. The new generation of scientists also has at their disposal a plethora of online resources offered by universities and private companies that focus specifically on AI. One example is Massive Open Online Courses (MOOCs), which are emerging as an affordable and popular option for those who want to deepen their knowledge of AI, from introductory courses to others on cutting-edge algorithms and advanced applications.
@@ 2.1.3 Computational resources
While AI is commonly perceived as an intangible technical system, it is de facto rooted in physical infrastructure and hardware. Yet, the role of physical assets and their associated computing capabilities – also known as AI compute – have been largely overlooked in policy circles and scholarly literature.
@@
Here, we measure the availability of AI compute by the presence of an HPC cluster within the focal researcher’s organization (although it should be noted that such compute infrastructure can also be used for non-AI workloads such as mathematical modeling and simulations). In very general terms, HPC is a technology that uses clusters of powerful processors, working in parallel, to process data and solve complex problems at high speeds (OECD, 2023). Unlike standard computing systems, HPC systems can handle multiple tasks simultaneously across multiple computer servers or processors with a centralized scheduler that manages the computing workload. The high cost of HPC can put this technology out of reach for most organizations, resulting in a significant “compute divide” within and between countries and institutions, as well as between the private sector and academia (Ahmed and Wahed, 2020; Ahmed et al., 2023). This is especially true for AI applications in some data-intensive scientific fields such as genomics and bioinformatics or particle physics where ML training and inferencing can be highly demanding in terms of memory and computational resources. The existence of a computational divide can therefore impede the adoption of AI and generate disparities in the productivity gains that AI can offer to science.
In summary, our hypothesis is that:
@@ 2.2.2 Taste for exploration
Exploration is intimately linked to curiosity, a personal trait that prompts individuals to explore uncharted territories. We think that curiosity is a useful construct for understanding scientists’ behavior in terms of technology adoption. Although psychologists have not reached a consensus on its definition, it is generally accepted that curiosity involves an intrinsic motivated desire for new information – an “appetite for knowledge”, or more formally “a form of cognitively induced deprivation that arises from the perception of a gap in knowledge or understanding(Loewenstein, 1994, p. 75).
@@ 3.1 Data
We use the entire OpenAlex database, which had more than 230 million scientific articles as of August 2022. OpenAlex provides us with information relevant to our study, namely titles and abstracts. Authors are disambiguated so that we can track the paper trail of our focal scientists and their co-authors. Thus, we can measure the evolution of the network of co-authors and the (publication) experience of the scientists who are part of it. Author affiliations are also clean and geographically localized. We add additional information about the organizations in our sample, particularly the university’s Shanghai ranking and the availability of high-performance computing (more on this below). Finally, we rely on a scientific journal categorization system provided by OpenAlex, called ‘concept’. In particular, we denote the scientific field of a paper by 0-level concept assigned to the journal of the paper (0-level concepts include broad areas such medicine, physics, biology, chemistry, etc.)
@@
The development of AI as a research tool is a relatively recent occurrence, and the widespread adoption of deep learning applications gained momentum in the early 2010s (Bianchini et al., 2022). More specifically, 2012 can be considered the beginning of the AI revolution, with significant advances in neural network technology (Krizhevsky et al., 2017; Sevilla et al., 2022). This prompts us to restrict attention to the first AI use in the period from 2012 (the year when AI took off) to 2020 (end of sample period). We further limit the analysis to researchers who started research after 1980; excluding older scientists at the end of their careers and, hence, in general less susceptible to adopt AI in research. Our sampling definition of focal scientists is concisely described in Figure 1.
@@ 3.3 Econometric strategy
where $p(y_{i,t}=1)$ is the probability that a paper produced applies AI conditional on a paper produced at all, $\phi$ denotes the logit-function. $\mathbf{h}_{i,t-1}$ is the log-transform of our measurements $\mathbf{H}_{i,t-1}$. In principle, some relevant capital and/or individual specific tendencies may be also unobserved $\nu_{i,t}$.With this estimation equation, the dynamics of AI technologies is effectively dealt with by introducing intercepts $\gamma_{s,t}$ that vary over time and scientific domains.
@@
These common factors can be removed in a conditional logit approach. First note that matching is on the outcome such that in all matches one individual $i$ adopts AI and the other scientist $j$ does not adopt AI $(y_{i,t}=1,y_{j,t}=0)$, or vice versa. The conditional logit model, takes into account that only two possible outcomes are possible, and we estimate the probability of one of them, where common factors have been pulled out of the second equation:
@@ 4.1 Main trends on AI adoption in science
A growing number of players: The adoption of AI technology by domain scientists is on the rise, with over 20,000 individuals incorporating it for the first time in 2020 alone. This pattern aligns with recent research on the diffusion of AI in science, showing that the number of AI publications has surged nearly five-fold in the past decade, constituting around 5% of total scientific publications. Most of this research has gradually shifted from AI development to its application, which currently represents some 70% of scientific activity (Arranz et al., 2023). Similar patterns have been observed for AI-related patent applications in OECD and non-OECD countries (Parteka and Kordalska, 2023). Surprisingly, despite technological advancements and increased resources for AI, the percentage of first-time users reusing technology remains low and stable at around 54% (13,211 re-use out of 23,918 first-use).
@@ 4.2.1 First-time AI adoption
Taken at face value the estimated coefficients suggest that being affiliated with an institution specialized in AI (AI inst. spe.=1), rather than not (AI inst. spe.=0), increases the odds-ratio for the focal scientist adopting AI compared to the matched scientist by 80%. [^7] To appreciate this point, note that our conditional logit model estimated on matched pairs allows for exactly two outcomes: either the focal scientist adopts AI and the matched scientist does not, or vice versa. The odds-ratio of the focal scientist adopting AI over the matched scientist adopting AI is $exp(\beta h_{i})/exp(\beta h_{j})$. A change of $i$’s factor from $h_{i}$ to $h^{\prime}_{i}$ thus implies a relative change of the odds-ratio by $exp(\beta(h^{\prime}_{i}-h_{i}))$. A binary factor changing from 0 to 1 yields a relative change of $exp(\beta)$, and a continuous factor (in logs in the model) being higher by 10% is about $exp(\beta\times 1.1)$. Hence, interpretation of coefficient estimates is easily facilitated by considering the exponential of the coefficient estimate. Scientists of more ‘excellent’ institutions with 10% higher citation rates have 30% higher odds-ratios of AI adoption.